ADNAN MUHAMMED

Grade 12,

a ball is left from the top of the building.a window which length is 3 meter is in the building the ball is pass the window in 0.5 second . if the velocity of the ball at the top of the window is Vt and bottom Vb( g=9.8) then which option is correct= 1.Vt+Vb=12 meter per second 2.Vt-Vb=4.9 meter per second 3.Vb.Vt=1 meter per second

a ball is left from the top of the building.a window which length is 3 meter is in the building the ball is pass the window in 0.5 second . if the velocity of the ball at the top of the window is Vt and bottom Vb( g=9.8) then which option is correct=
1.Vt+Vb=12 meter per second
2.Vt-Vb=4.9 meter per second 
3.Vb.Vt=1 meter per second

Grade:12

2 Answers

Dhruvit Raithatha
6 years ago
\\ s = 3\ ms^{-1}, t = 0.5\ s,g = 9.8 ms^{-1} \\ s = v{_t}t + \frac{1}{2}at^2 \\ 3 = \frac{v{_t}}{2} +\frac{9.8}{8} \\ v{_t} = \frac{24 - 9.8}{4} = \frac{14.2}{4} \approx 3.5 \\ v{_b} = v{_t} + gt = 3.5 + (9.8)(0.5) \approx 8.5 \\ \therefore\text{ (a) and (b) are correct}\\ s = 3\ ms^{-1}, t = 0.5\ s,g = 9.8 ms^{-1} \\ s = v{_t}t + \frac{1}{2}at^2 \\ 3 = \frac{v{_t}}{2} +\frac{9.8}{8} \\ v{_t} = \frac{24 - 9.8}{4} = \frac{14.2}{4} \approx 3.5 \\ v{_b} = v{_t} + gt = 3.5 + (9.8)(0.5) \approx 8.5 \\ \therefore\text{ (a) is correct}
Dhruvit Raithatha
6 years ago
Correction:
\text{Ignore the first image. Only (a) is correct, as } v{_t} - v{_b} = - 9.8
Donotreadthisitsjustaplaceholdertofulfilltheminimumhundredcharachterrulegod whyareyoudtillreadingthisitoldyouitsuseless.